### v3

# The 2D+ Dynamic Articulatory Model DYNARTmo: Tongue-Palate Contact Area Estimation

Abstract
We present DYNARTmo, a dynamic articulatory model designed to visualize speech articulation processes in a two-dimensional midsagittal plane. The model builds upon the UK-DYNAMO framework and integrates principles of articulatory underspecification, segmental and gestural control, and coarticulation. DYNARTmo simulates six key articulators based on ten continuous and six discrete control parameters, allowing for the generation of both vocalic and consonantal articulatory configurations. The current implementation is embedded in a web-based application (SpeechArticulationTrainer) that includes sagittal, glottal, and palatal views, making it suitable for use in phonetics education and speech therapy. While this paper focuses on the static modeling aspects, future work will address dynamic movement generation and integration with articulatory-acoustic modules.
1. Introduction
Articulatory models describe the stage of speech where neuromuscular activation patterns lead to movement patterns of speech articulators (i.e., lips, tongue, velum, lower jaw) with the goal of generating audible speech signals as part of speech communication processes. The motivation for developing articulatory models is manifold. (i) Models unfold the process of speech articulation and acoustic signal generation and thus help to understand the basic mechanisms of articulation and coarticulation. (ii) If coupled with an articulatory-acoustic module, an articulatory model provides insight into the mechanisms of acoustic signal generation at the vocal folds and within the vocal tract. (iii) Articulatory models may serve in the future as high-quality speech synthesizers, capable of generating different voices (i.e., different speakers), different emotional connotations, and more—beyond merely transmitting a spoken message. However, corpus-based synthesis systems currently produce the most natural speech quality, and many research questions remain to be addressed before high-quality articulatory synthesizers can be realized (Campbell,, 2005; Kröger,, 2022).
At present, articulatory models are mainly used as research tools to study articulation and coarticulation or, when combined with articulatory-acoustic modules, to investigate aspects of acoustic signal generation, including sublaryngeal, laryngeal, and supralaryngeal aerodynamics (Birkholz,, 2013; Kröger,, 2022; Fan et al.,, 2024). Other models are used in speech therapy to visually present articulator positions ((Kröger et al.,, 2005)) or articulatory contact information (e.g., tongue–palate or labial contact), offering somatosensory feedback to help improve our imagination of articulation. Examples include electropalatography (EPG), which provides real‑time visualization of tongue–palate contact during speech (e.g. (Hardcastle et al.,, 1991; Nordberg et al.,, 2011)), and ultrasound visual biofeedback (U‑VBF), displaying dynamic tongue surface movements in therapy for speech sound disorders ((Sugden et al.,, 2019; Preston et al.,, 2017)). Moreover, articulatory models can be used to simulate speech learning or relearning processes in therapeutic contexts, or as part of second and third language learning. In that case, articulatory models need to be integrated into large-scale neurobiologically motivated neural models of language processing (Kröger,, 2023).
Many aspects of speech articulation can be represented in the midsagittal plane. Therefore, early articulatory models were typically two-dimensional (Henke,, 1966; Mermelstein,, 1973; Coker,, 1976; Maeda,, 1979; Heike,, 1979). These models primarily aimed at generating high-quality synthetic speech. However, by the end of the 20th century, corpus-based synthesizers such as the Klatt synthesizer (Klatt,, 1980) and newer systems like corpus-bases speech synthesizers (Campbell,, 2005) had outperformed articulatory models in speech quality. Consequently, articulatory models—including articulatory-acoustic variants—became more focused on their use as research tools, particularly in studying speech acquisition, articulation, and coarticulation across languages and speaker groups. As a result, increasingly detailed models emerged that incorporate 3D spatial information, neuromuscular control schemes, and biomechanical modeling of muscles and articulatory tissues (Kröger,, 2022).
The model described in this paper, DYNARTmo (DYNamic ARTiculatory model), is a further development of the UK-DYNAMO (Universität Köln-DYNamic Articulatory MOdel) originally introduced by Heike (Heike,, 1979). The UK-DYNAMO was inspired by Henke’s approach (Henke,, 1966) and later used in a computer-based tool for speech training (Heike,, 1989). Like many models of its time, UK-DYNAMO was a 2D system based on midsagittal contours of articulator surfaces, specifically German vowel and consonant contours from Wängler (Wängler,, 1958). It employed temporal interpolation and a basic form of coarticulation as suggested by Henke.
DYNARTmo was initially introduced by Kröger (Kröger,, 1992), who incorporated the concept of articulatory underspecification to derive minimal rules for articulation and coarticulation. This approach was elaborated through two control concepts: a gestural control scheme (Kröger,, 1993; Kröger et al.,, 1995) and a segmental approach (Kröger,, 2001). Starting in the 2000s, DYNARTmo was updated with geometries derived from MRI data (Kröger et al.,, 2000, 2004), offering a more realistic articulatory basis. Later, it was embedded into a large-scale neurobiological speech production model (Kröger et al.,, 2014), incorporating concepts such as primary and secondary articulators, muscle bundles, and neuromuscular activation patterns (Kröger and Bekolay,, 2022).
This paper presents the current implementation of DYNARTmo as part of the SpeechArticulationTrainer web application, available at https://speechtrainer.eu/projects.htm. In addition to the midsagittal view, the app now includes glottal (superior) and palatal (inferior) views. Together, these three perspectives offer an integrated spatial understanding of articulatory processes. While the focus of this paper is on the static aspects of the model, future work will address the generation of articulatory movement patterns.
2. DYNARTmo: Static Aspects
DYNARTmo simulates six model articulators—lips, tongue tip, tongue dorsum, lower jaw, velum, and glottis—which are controlled by a set of ten continuously scaled control parameters (see Table 1 and Figure 1). These control parameters govern the displacement of each articulator from its neutral or rest position. Functionally, the parameters are grouped into three vocalic parameters, three consonantal parameters, one parameter for velum control, and three parameters for phonatory control. It is important to note at this point that the lower jaw is not directly controlled by any of the parameters introduced here (see Discussion).
Figure 1: Midsagittal views generated by DYNARTmo. (a) shows the neutral configuration of all articulator contours for a central vocalic state. (b–h) display paired overlays of articulator contours, illustrating edge positions (displacements) for individual control parameters: (b) vocalic height (high–low) for a front vowel position; (c) vocalic front–back for a high vowel position; (d) lip rounding (spread–rounded) for a high front vowel; (e) labial aperture from vocalic to full closure (for a low vowel); (f) tongue tip height from vocalic to closure (for a low vowel); (g) tongue dorsum height from vocalic to closure (for a low vowel); (h) velum height (velopharyngeal aperture: closed–open) for the neutral vocalic position.
The vocalic control parameters specify tongue height (high–low), tongue position along the front–back axis, and the degree of lip rounding (spread–rounded). The consonantal parameters define the degree of constriction generated by the lips, tongue tip, or tongue dorsum. These values are relative to the current vocalic articulator position and describe, for example, the elevation of the tongue tip or dorsum toward closure, or respectively the reduction of labial aperture from an open vocalic state to full closure.
In addition, further control parameters are defined for the velopharyngeal and laryngeal–sublaryngeal systems (Table 1). Velum height controls the opening of the velopharyngeal port. Glottal aperture sets the position of the arytenoid cartilages, thereby determining the resting distance of the vocal folds for breathing, voiceless sounds, or phonation. Vocal fold tension controls the pitch (i.e., the fundamental frequency) of phonation. Lung pressure governs the production of airflow required for phonation and determines the loudness (intensity) of the resulting speech signal.
Table 1: Continuously scaled control parameters used in DYNARTmo, including affected articulators and their functional roles. Vocalic parameters define global vocal tract shaping; consonantal parameters govern localized constrictions. Additional parameters control velopharyngeal aperture, glottal configuration, and subglottal pressure. (Speech sounds are transcribed using SAMPA; see (Wells,, 1997))
Type Name Affected Articulators Function
voc vocalic high–low tongue body, lower jaw vertical shaping of the vocal tract (vowel height)
voc vocalic front–back tongue body, lower jaw horizontal shaping of the vocal tract (vowel position)
voc lips spread–round lips lip rounding, relevant for vowels and some consonants (e.g., /S/)
cons labial aperture lips, lower jaw constriction for labial consonants (e.g., /b/, /p/, /m/)
cons tongue tip height tongue tip, tongue body, lower jaw constriction in the apical region (e.g., /d/, /t/, /n/)
cons tongue dorsum height tongue dorsum, tongue body, lower jaw constriction in the dorsal region (e.g., /g/, /k/, /N/)
vel velum height velum degree of velopharyngeal opening (nasality control)
glott glottal aperture arytenoid cartilages (larynx) voicing control: open for voiceless, closed for phonation
glott vocal fold tension vocal folds (larynx) pitch (fundamental frequency) control via tension modulation
lung lung pressure sublaryngeal system controls airflow intensity and loudness of the speech signal
In addition to the ten continuously scaled control parameters (see Table 1), the model includes six discrete parameters that define the place and manner of consonantal constrictions (see Table 2). The value range of the continuous parameters varies depending on their functional category: for vocalic parameters, values range from 1-1 to $1$, where 0 indicates the neutral articulator position. The same 1-1 to $1$ scale is used for velopharyngeal and glottal aperture, with $1$ indicating full opening, 0 indicating closure, and 1-1 representing tight closure. For parameters governing consonantal constrictions (i.e., labial, apical, and dorsal), as well as for vocal fold tension and lung pressure, the scale ranges from 0 to $1$. A value of 0 indicates no active constriction (i.e., the articulator remains in its current vocalic position), while a value of $1$ represents full articulatory closure or maximum effort.
The discrete control parameters (Table 2) define both the place and manner of consonantal articulation. The manner of articulation distinguishes between: (i) full closure (as in plosives and nasals), (ii) fricative closure, in which the airstream is focused through a midsagittal groove (or, in labiodental fricatives, between the lower teeth and upper lip), and (iii) lateral closure, which involves strong central raising and simultaneous lateral lowering of the tongue surface, as in laterals. These manners correspond to distinct articulatory configurations or shapes.
The place of articulation is categorized according to anatomical regions: bilabial or labiodental for the lips; dental, alveolar, or post-alveolar for the tongue tip; and palatal or velar for the tongue dorsum. It should be noted that although tongue tip and tongue dorsum positions are represented by continuously scaled parameters in the current model implementation, they are typically assigned discrete values to reflect linguistically relevant articulatory targets during consonant production.
Table 2: Discrete control parameters in DYNARTmo, defining place and manner of consonantal constriction. Each articulator is associated with specific labels for place and manner of articulation. Manner defines the constriction shape; place defines the anatomical target region.
Articulator Parameter Type Possible Labels
lips place bilabial, labiodental
lips manner full (vocalic), near (central groove or lower teeth row)
tongue tip place dental, alveolar, post-alveolar
tongue tip manner full (vocalic), near (central groove), lateral (lowering)
tongue dorsum place palatal, velar
tongue dorsum manner full (vocalic), near (central groove)
The set of articulatory control parameters does not directly determine the positions of the model articulators. Rather, these parameters serve a functional role: they specify consonantal articulation in terms of the degree and location of vocal tract constrictions, and vocalic articulation in terms of canonical phonetic-linguistic categories such as high–low, front–back, and spread–rounded (cf.(IPA,, 1999): International Phonetic Association).
The actual specification of articulator positions is carried out in a subsequent computational step, based on the values of all control parameters at a given time point. This step involves computing the vertical and horizontal positions of the tongue tip, tongue dorsum, and lips, as well as the vertical positions of the velum, lower jaw, and larynx (see Fig. 1). Conceptually, this transformation can be understood as the mapping of high-level vocal tract parameters—such as constriction degree and location for consonants, or vocal tract tube shape for vowels—onto concrete articulator geometries, akin to the task dynamics approach proposed by (Saltzman and Munhall,, 1989).
The calculation of model articulator positions follows a four-step procedure: (i) computation of vocalic articulator positions as a baseline, (ii) computation of primary articulator positions for consonantal constrictions, (iii) computation of the lower jaw position as a mediating articulator for labial, apical, or dorsal constriction formation, and (iv) adjustment of tongue and lip positions in response to the jaw movement.
This stepwise approach reflects the inherent parallelism of vocalic and consonantal articulation in fluent speech. During vocalic segments, the articulators are often still moving away from a previous consonantal constriction or already transitioning toward the next one. Similarly, consonants are never produced in isolation but occur in vocalic context, which means that the vocal tract shape during constriction formation is inevitably influenced by both preceding and following vowels. In this sense, aspects of dynamic articulation are already embedded within the otherwise static modeling framework proposed in this paper.
Step 1a: Vocalic articulator positions for the tongue, jaw, velum, larynx, and lips are interpolated based on the high–low and front–back parameters. This interpolation is derived from three prototypical vowel shapes corresponding to the cardinal vowels /i/, /a/, and /u/. In the neutral case (high–low = 0, front–back = 0), an intermediate position is computed (see Fig. 1a). For edge cases, high–low = 1 or –1 represents a high or low vowel respectively (Fig. 1b), and front–back = 1 or –1 indicates a front or back vowel respectively (Fig. 1c).
Step 1b: The interpolated lip shape is further refined by incorporating a fourth contour: the lips and jaw configuration for the rounded front vowel /y/. The spread–rounded parameter controls this interpolation, with values of 0 (spread) and 1 (rounded) corresponding to Fig. 1d.
Step 2: Based on the current values of the consonantal constriction parameters—i.e., labial aperture, tongue tip height, and tongue dorsum height—the model interpolates between the vocalic articulator contour (consonantal parameter = 0) and the full constriction or closure (consonantal parameter = 1). This applies to lips, tongue tip, and tongue dorsum (Fig. 1e–g).
Step 3: The lower jaw position is then interpolated between the value determined by the vocalic state (determined by vocalic control parameters) and the consonantal constriction parameter value determining the degree of consonantal constriction. In many cases, this results in a further elevation of the jaw beyond its basic vocalic position to support a consonantal target (Fig. 1e–g).
Step 4: This jaw-induced elevation (Step 3) in turn leads to secondary adjustments in other articulators: (i) In the case of labial constrictions (Fig. 1e), the tongue tip is elevated; (ii) for dorsal constrictions (Fig. 1g), the lower lip is elevated; (iii) for apical constrictions (Fig. 1f), both the tongue dorsum and lips are affected by the jaw’s vertical shift.
3. Integration of DYNARTmo into a Web App for Speech Therapy and Language Learning
One of the current objectives is to use the model as a visual feedback tool for learning or re-learning speech sound production in therapeutic contexts (see Fig. 2). Individuals affected by speech disorders are often highly motivated to engage with tools that offer visual support to improve their articulatory capabilities.
Figure 2: Screenshot from the SpeechArticulationTrainer web application, showing the sagittal, glottal, and palatal views for four speech sounds: the vowels [a] and [i], and the consonants [t] and [k]. (The tongue-palate contact pattern for /t/ lacks lateral tongue contact; see Discussion.)
A second goal is to integrate the model into phonetics instruction, for example as part of course curricula in linguistics or speech-language pathology (logopaedics).
In addition to displaying the form and displacement of articulators from their rest positions in a sagittal plane view (see Fig. 1), the SpeechArticulationTrainer web application also provides: (i) a glottal view (transverse plane, superior perspective), and (ii) a palatal view (transverse plane, inferior perspective), as illustrated in Figure 2.
The glottal view provides visual feedback on the state of the vocal folds: abducted for phonation (voiced sounds) versus adducted for the production of voiceless sounds. The palatal view illustrates tongue–palate contact regions, which correspond to tactile sensations perceived by the speaker during articulation.
Arrows (and dots) indicate the airflow direction, while a grey dot marks the presence of subglottal pressure, which is necessary for sound generation.
It should be noted that plosive consonants cannot be fully captured in static images like those shown in Figure 2. Static frames do not convey the release phase of the constriction or the subsequent noise burst characteristic of plosives. For this reason, the application also includes the ability to generate animated sequences of articulatory movement patterns—at the syllable, word, or short phrase level—enabling more dynamic and comprehensive visualization.
4. Discussion
One central goal of this study is to propose a simplified articulatory model that operates without the full task-dynamics formalism, while still reflecting key distinctions commonly found in more complex frameworks such as the Task Dynamics Approach ((Saltzman and Munhall,, 1989)). Notably, our model implicitly includes a distinction between articulator space and tract variable space. However, this distinction is not strictly separated in our formulation. For instance, articulatory parameters such as tongue dorsum height or velum position are directly mapped to constriction properties, thereby bypassing an explicit intermediate representation of tract variables. This deliberate simplification serves to make the model computationally tractable and visually transparent, especially for educational or diagnostic applications. Nonetheless, the blending of articulator-based and functionally-defined parameters might lead to a certain conceptual ”blurring” that deserves clarification. Future refinements of the model may aim to more clearly disentangle these parameter spaces, or at least highlight their interplay more explicitly, in order to facilitate theoretical comparisons with established models of speech production.
As outlined in Chapter 2, our model differentiates between vocalic and consonantal articulation. Moreover, we emphasize there that these two articulatory domains often temporally overlap during fluent speech. This overlap is reflected in our approach by describing the interplay between vocalic and consonantal gestures within a static articulatory model. This conceptual treatment of vowel–consonant interaction can thus be seen as a first step toward modeling temporal coordination within articulation. In that sense, these considerations already incorporate elements of dynamic articulation into the otherwise static modeling framework presented in this paper.
It should be noted that the tongue-palate contact patterns presented in this paper are based on a preliminary version of our 3D module for estimating tongue-palate contact. As shown in Fig. 2, the current model fails to represent lateral tongue elevation at the molar region in the case of the alveolar plosive /t/, resulting in an incorrect apical-only contact. A more complete 3D implementation —including detailed modeling of the tongue groove—will be incorporated in forthcoming versions of the model, covering both plosive and fricative contexts.
5. Further Work
Several extensions of the current model are planned for future development and publication. These include: (i) the implementation of a virtual 3D model for generating detailed lingual–palatal contact patterns, (ii) a comprehensive description of segmental and gestural control mechanisms for generating dynamic articulatory movement patterns, and (iii) the integration of aerodynamic representation within both the model and the SpeechArticulationTrainer web application.
In addition, we aim to reintegrate an articulatory–acoustic module into a neural model framework. This will allow DYNARTmo to serve as a component within a large-scale, neurobiologically plausible model of speech production and perception (Kröger,, 2023).
The source code of the current implementation of DYNARTmo is included as Zip-File.
Acknowledgements
I am grateful to Georg Heike for introducing me to the fascinating field of articulatory modeling. I also wish to thank Catherine Browman, Louis Goldstein, Elliot Saltzman, and Philip Rubin for their inspiring discussions during my visit to Haskins Laboratories in New Haven, CT (1994), as well as during a phonology conference in Graz, Austria (1992) with Catherine Browman. These interactions had a formative influence on the early development of the present model.
Supplementary Material
The Python source code used to implement the static visualization routines of DYNARTmo (static model part) is provided as a supplementary ZIP archive (dynArtMo.zip) attached to this arXiv submission. The archive contains the Jupyter notebook articulatoryModel.ipynb, which includes all relevant computational routines for generating the sagittal view of vocal tract articulators, as well as a subfolder data_sagi containing the contour data for the reference vocal tract shapes. This material allows full reproducibility of the visualizations presented in the main text.